A spinner for a child's game has the numbers 1 through

9 evenly spaced. If a child spins, find each probability:

a. The number is divisible by 3.

b. the number is greater than 7.

c. The number is an even number.


1
Expert's answer
2022-02-15T08:28:45-0500

Sample space = {1,2,3,4,5,6,7,8,9}

Total number of outcomes = 9

a. Number divisible by 3 in the sample space = 3, 6, 9

Thus, no. of favourable outcomes = 3

Required probability = "\frac{3}{9}=\frac{1}{3}"

b. Number greater than 7 in the sample space = 8, 9

Thus, no. of favourable outcomes = 2

Required probability = "\frac{2}{9}"

c. Even number in the sample space = 2, 4, 6, 8.

Thus, no. of favourable outcomes = 4

Required probability = "\frac{4}{9}"